OpenStudy (anonymous):

A line passes through (-3,-1) slope of -5

OpenStudy (jojo4eva):

y=-5x-16

OpenStudy (jojo4eva):

use the same thing as last time use the equation y-y1=m(x-x1)

OpenStudy (the_fizicx99):

Y-y1 = m(x-x1) y-(-1) = -5(x-(-3)) y + 1 = -5(x + 3) y + 1 = -5x - 15 -1 -1 y = -5x - 16

OpenStudy (anonymous):

hmm okay that helps alot

OpenStudy (anonymous):

A line passes through the point (4,8) and has a slope of .6

OpenStudy (anonymous):

soo y-(8) = 6(x-(4))

OpenStudy (jojo4eva):

exactly and do the work from there

OpenStudy (jojo4eva):

and u get y=6x-16

OpenStudy (anonymous):

hmm okay i got it now

OpenStudy (jojo4eva):

kk good juss lemme kno if you have any other problems

OpenStudy (anonymous):

okay!